# On the tractability of SHAP explanations under Markovian distributions

### v1

@@ 1 Introduction
Since its introduction in the seminal paper (Lundberg and Lee, 2017), the local explanatory (SHapley Additive exPlanations) SHAP method gained increasing popularity in the field of interpretable ML. Nevertheless, one of its main limitations pertains to its computational intractabilty: In general, computing the SHAP score is NP-Hard (Bertossi et al., 2020; den Broeck et al., 2021). A recent body of works showed promising positive results regarding the tractability of computing the SHAP score under particular configurations. In (Lundberg et al., 2020), a polynomial-time algorithm to compute exactly the SHAP score for the family of decision trees, commonly known as TreeSHAP, has been proposed. A more rigorous proof for the tractability of various families of boolean functions has been provided in (den Broeck et al., 2021). (Arenas et al., 2023) extended these positive complexity results to cover the family of Decomposable Deterministic circuits which includes the family of decision trees among other classes of boolean circuits.
@@
All these reported results in the literature are, however, derived under the feature independence assumption. Although practical for its simplicity, this assumption is often irrealistic in real-case scenarios. A slight relaxation of this assumption has been examined in (den Broeck et al., 2021) through the lens of complexity theory by considering the family of naïve bayes models and empirical distributions. Computing the SHAP score for the family of decision trees under this relaxed assumption has been proven to be #P-Hard in both these settings.
@@
Between independent distributions and latent variable models, an intermediate class of distributions that hasn’t been explored yet is the class of Markovian distributions. Markovian distributions constitute an interesting class of distributions that incorporate a degree of feature correlation often considered sufficient to model various stochastic phenomena (Bassler, Gunaratne, and McCauley, 2006; Kampen, 2007; Goutsias and Jenkinson, 2013).
@@
Previous works examining the complexity of computing the SHAP score were mostly directed towards families of boolean functions. In this article, we shift our focus to sequential models, in particular the family of weighted automata (WAs). WAs offer a powerful formalism for modeling sequential tasks and encompass a large family of classical models, including Deterministic and Non deterministic finite automata, Hidden Markov Models, and has been shown to be equivalent to second-order linear RNNs (Rabusseau, Li, and Precup, 2019). They have been employed in various applications, such as NLP (Knight and May, 2009), speech processing (Pereira and Riley, 1996; Mohri, Pereira, and Riley, 2008) and image processing (Culik and Kari, 1993).
@@
Recently, a line of works proposed WAs as proxy interpretation models for neural models (Okudono et al., 2020; Eyraud and Ayache, 2021; Weiss, Goldberg, and Yahav, 2019; Lacroce, Panangaden, and Rabusseau, 2021). All these works are motivated by the implicit assumption that WAs enjoy better transparency than their neural counterparts. However, the existing litterature lacks a formal argument to substantiate this claim. One of the primary motivations of this work is to shed some light on this issue.
$\bullet$ Contributions. The main results presented in this article are given as follows:
@@ 2 Background
    The product operator: The product operator, also known as the hadamard product (Droste and Gastin, 2009; Mohri, 2004), takes two languages $f,~{}g$ over $\Sigma^{*}$ and outputs the product language $f\cdot g$. We shall employ the notation $f\otimes g$ to refer to the product language of $f$ and $g$.
@@ 2.1 Weighted Automata/Transducers
((Denis and Esposito, 2008)) Let $\Sigma$ be an alphabet and $n>0$ be an integer. A WA $A$ over $\Sigma^{*}$ is represented by a tuple $<\alpha,\{A_{\sigma}\}_{\sigma\in\Sigma},\beta>$ where $A_{\sigma}\in\mathbb{R}^{n\times n}$ is the transition matrix associated to a symbol $\sigma$ in $\Sigma$, and $\alpha$ (resp. $\beta$) are vectors in $\mathbb{R}^{n}$ that represent the initial (resp. final) vectors. The integer $n$ is called the size of $A$, denoted $\texttt{size}(A)$.
@@ A WA $A=<\alpha,\{A_{\sigma}\}_{\sigma\in\Sigma},\beta>$ over $\Sigma^{*}$ computes the language
$\bullet$ Weighted transducers. Weighted transducers (WTs) represent the analogous version of WAs adapted to model seq2seq languages. It has been employed in applications including speech processing (Mohri, Pereira, and Riley, 2008; Lehr and Shafran, 2010), machine translation (Kumar, Deng, and Byrne, 2006) and image processing (Culik and Friš, 1995)
@@
where $(w,u)\in\Sigma^{*}\times\Delta^{*}$ such that $|w|=|u|$  [^2] Note that WTs described via linear representations assign an output value only to a pair of sequences that have the same length. This formalism is sufficient in the context..
@@
Stemming its root from the field of cooperative game theory (Deng and Papadimitriou, 1994), the SHAP framework is built on top of an analogy between cooperative games and the local explainability problem of ML models. A cooperative game is described by a set of players $N$ and a value function $v$ that assigns a generated wealth for each subset of players, referred to as a coalition, cooperating in the game. By analogy, in the context of explainable ML, the players are the input features of a ML model subject to explanatory analysis. And, the value assigned to a coalition is equal to the the expected model’s output conditioned on the event that the features forming the coalition possess a value equal to the instance to explain.
@@
Formally, let $M$ be a model that computes a function $f_{M}$ from a discrete set $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$ to $\mathbb{R}$, and $P$ be a probability distribution over $\mathcal{X}$. For an input $x\in\mathcal{X}$, and an integer $i\in[n]$, the SHAP score assigned to the $i$-th feature for the instance $x$ is given as (Lundberg and Lee, 2017):
@@
    $\forall(w^{\prime},p)\in\Sigma^{|w|}\times\Sigma_{\#}^{|w|}:~{}~{}g_{w,i,P}^{(2)}(w^{\prime},p)=P(w^{\prime}|w^{\prime}\in L_{\texttt{swap}(p,i)})$ (11)
@@
Expressions (8) and (9) reduce the problem of computing $\text{SHAP}_{1}$ and $\text{SHAP}_{2}$ to that of performing operations over a language $f_{w,k}$ and two seq2seq languages, $g_{w,P}^{(1)},~{}g_{w,i,P}^{(2)}$ whose properties are given by equations (10) and (11) , respectively. The missing link to complete the proof of lemma 3.3 is to prove that a WA that implements the language $f_{w,k}$, and WTs that compute seq2seq languages $g_{w,k}^{1},~{}g_{w,k}^{2}$ whose properties are given in lemma 3.4 can be constructed in polynomial time.
3.2.2 Step 2.b: Construction of WAs/WTs that compute $f_{w,k},~{}g_{w,P}^{1},~{}g_{w,i,P}^{2}$
@@
It’s worth noting that since $\mathcal{A}_{1},~{}\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ run in time polynomial in their respective input instances implies that the size of their output machines is also polynomial in the size of their input instance [^3] FP $\subset$ FPSPACE. This fact will appear explicitly in the constructive proof of lemma 3.5.
@@
$\bullet$ Disjoint-DNFs. A d-DNF is a boolean formula formed as a disjunction of conjunctive clauses that are pairwise mutually exclusive.
@@
A Markovian distribution $P$ over a boolean random vector of dimension $N$ is given as:  [^4] In reality, the set of Markovian distributions is more general and ranges over all possible permutations of $[N]$. For the sake of simplification, we consider only the case of Markovian distributions under the identity permutation. We note that this assumption doesn’t have an impact on the tractability results presented in this section.
@@
This result follows immediately from theorem 4.1, and the fact that given an arbitrary decision tree in DT, an equivalent d-DNF can be constructed in polynomial time with respect to the size of the decision tree (Property 1, (Aizenstein and Pitt, 1992)). ∎
@@ 5 Conclusion
$A_{w,P}$ refers to the WA that computes the language expressed in 18.
Given a string $w\in\Sigma^{*}$ and $P\in\texttt{MARKOV}$. A Markvovian distribution over the finite support $\Sigma^{|w|}$ can be easily simulated by a DWA. The construction consists at maintaining in the state memory of the DWA the position reached so far in the sequence and the last generated symbol. These two pieces of information are sufficient to simulate a Markovian distribution.
@@
For an input instance $<\sigma^{\prime},n,\sigma,m,P>$, the quantity $\mathbf{G}(\sigma^{\prime},n,\sigma,m,P)$ refers to the conditional probability of generating a symbol in $L_{\sigma^{\prime}}$ at position $m$ given that the symbol $\sigma$ has been generated at position $n$. In essence, the computational problem $\mathbf{G}$ is reduced to the classical problem of inference in Bayesian Networks (Koller and Friedman, 2009). In general, the exact inference in Bayesian Networks is intractable (Tacettin and Ünlüyurt, 2005). However, in our case, leveraging the Markovian structure of the probability distribution enables building a tractable solution for the problem using a dynamic programming approach.

### v2

@@ 1 Introduction
Since its introduction in the seminal paper (Lundberg & Lee, 2017), the local explanatory (SHapley Additive exPlanations) SHAP method gained increasing popularity in the field of interpretable ML. Nevertheless, one of its main limitations pertains to its computational intractabilty: In general, computing the SHAP score is NP-Hard (Bertossi et al., 2020; den Broeck et al., 2021). Recent studies have shown positive results regarding the tractability of computing the SHAP score under specific configurations. In particular, (Lundberg et al., 2020) proposed a polynomial-time algorithm, known as TreeSHAP, that purported to compute exactly the SHAP score for tree-based models. However, subsequent research (den Broeck et al., 2021; Arenas et al., 2023) has identified flaws in the main claim of TreeSHAP. Indeed, other works have demonstrated that the TreeSHAP algorithm is an implementation of interventional SHAP, another variant of the SHAP score (Janzing et al., 2020). A more rigorous proof for the tractability of the original SHAP score for various families of boolean functions has been provided in (den Broeck et al., 2021) when features are assumed to be independent. (Arenas et al., 2023) extended these positive results to cover the family of Decomposable Deterministic circuits which includes the family of decision trees among other classes of boolean circuits.
@@
All the tractability results reported in the literature are, however, derived under the feature independence assumption. Although practical for its simplicity, this assumption is often irrealistic in real-case scenarios. A slight relaxation of this assumption has been examined in (den Broeck et al., 2021) through the lens of complexity theory by considering the family of naïve bayes models and empirical distributions. Computing the SHAP score for the family of decision trees under this relaxed assumption has been proven to be #P-Hard in both these settings.
@@
Between independent distributions and latent variable models, an intermediate class of distributions that hasn’t been explored yet is the class of Markovian distributions. Markovian distributions constitute an interesting class of distributions that incorporate a degree of feature correlation often considered sufficient to model various stochastic phenomena (Bassler et al., 2006; Kampen, 2007; Goutsias & Jenkinson, 2013).
@@
Previous works examining the complexity of computing the SHAP score were mostly directed towards families of boolean functions. In this article, we shift our focus to sequential models, in particular the family of weighted automata (WAs). WAs offer a powerful formalism for modeling sequential tasks and encompass a large family of classical models, including Deterministic and Non deterministic finite automata, Hidden Markov Models, and has been shown to be equivalent to second-order linear RNNs (Rabusseau et al., 2019). They have been employed in various applications, such as NLP (Knight & May, 2009), speech processing (Pereira & Riley, 1996; Mohri et al., 2008) and image processing (Culik & Kari, 1993).
@@
Recently, a line of works proposed WAs as proxy interpretation models for neural models (Okudono et al., 2020; Eyraud & Ayache, 2021; Weiss et al., 2019; Lacroce et al., 2021). All these works are motivated by the implicit assumption that WAs enjoy better transparency than their neural counterparts. However, the existing litterature lacks a formal argument to substantiate this claim. One of the primary motivations of this work is to shed some light on this issue.
The work presented in this article will primarily address the original formulation of the SHAP score, as introduced by (Lundberg & Lee, 2017) in their seminal paper. This specific variant of the SHAP score has proven to be particularly challenging from a computational viewpoint, and extends, up to a distributional shift, other variants such as the baseline SHAP (Sundararajan & Najmi, 2020). It’s worth noting, however, that the axiomatic basis of the original SHAP score has been recurrently disputed in the academic discourse with several works discussing its limitation to capture elementary desirable properties of local models’ explanations (Janzing et al., 2020; Sundararajan & Najmi, 2020; Huang & Marques-Silva, 2023).
The main results presented in this article are given as follows:
@@ 2 Background
    The product operator: The product operator, also known as the hadamard product (Droste & Gastin, 2009; Mohri, 2004), takes two languages $f,~{}g$ over $\Sigma^{*}$ and outputs the product language $f\cdot g$. We shall employ the notation $f\otimes g$ to refer to the product language of $f$ and $g$.
@@ 2.1 Weighted Automata/Transducers
((Denis & Esposito, 2008)) Let $\Sigma$ be an alphabet and $n>0$ be an integer. A WA $A$ over $\Sigma^{*}$ is represented by a tuple $<\alpha,\{A_{\sigma}\}_{\sigma\in\Sigma},\beta>$ where $A_{\sigma}\in\mathbb{R}^{n\times n}$ is the transition matrix associated to a symbol $\sigma$ in $\Sigma$, and $\alpha$ (resp. $\beta$) are vectors in $\mathbb{R}^{n}$ that represent the initial (resp. final) vectors. The integer $n$ is called the size of $A$, denoted $\texttt{size}(A)$.
@@ A WA $A=<\alpha,\{A_{\sigma}\}_{\sigma\in\Sigma},\beta>$ over $\Sigma^{*}$ computes the language
$\bullet$ Weighted transducers. Weighted transducers (WTs) represent the analogous version of WAs adapted to model seq2seq languages. It has been employed in applications including speech processing (Mohri et al., 2008; Lehr & Shafran, 2010), machine translation (Kumar et al., 2006) and image processing (Culik & Friš, 1995)
@@
where $(w,u)\in\Sigma^{*}\times\Delta^{*}$ such that $|w|=|u|$.
@@
Stemming its root from the field of cooperative game theory (Deng & Papadimitriou, 1994), the SHAP framework is built on top of an analogy between cooperative games and the local explainability problem of ML models. A cooperative game is described by a set of players $N$ and a value function $v$ that assigns a generated wealth for each subset of players, referred to as a coalition, cooperating in the game. By analogy, in the context of explainable ML, the players are the input features of a ML model subject to explanatory analysis. And, the value assigned to a coalition is equal to the the expected model’s output conditioned on the event that the features forming the coalition possess a value equal to the instance to explain.
@@
Formally, let $M$ be a model that computes a function $f_{M}$ from a discrete set $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$ to $\mathbb{R}$, and $P$ be a probability distribution over $\mathcal{X}$. For an input $x\in\mathcal{X}$, and an integer $i\in[n]$, the original SHAP score assigned to the $i$-th feature for the instance $x$ is given as (Lundberg & Lee, 2017):
@@
    $\forall(w^{\prime},p)\in\Sigma^{|w|}\times\Sigma_{\#}^{|w|}:~{}~{}g_{w,i,P}^{(2)}(w^{\prime},p)=P(w^{\prime}|w^{\prime}\in L_{p^{\prime}})$ (11)
    where $p^{\prime}=\texttt{swap}(p,i)$
@@
Expressions (8) and (9) reduce the problem of computing $\text{SHAP}_{1}$ and $\text{SHAP}_{2}$ to that of performing operations over a language $f_{w,k}$ and two seq2seq languages, $g_{w,P}^{(1)},~{}g_{w,i,P}^{(2)}$ whose properties are given by equations (10) and (11) , respectively. The missing link to complete the proof of lemma 3.3 is to prove that a WA that implements the language $f_{w,k}$, and WTs that compute seq2seq languages $g_{w,P}^{(1)},~{}g_{w,i,P}^{(2)}$ whose properties are given in lemma 3.4 can be constructed in polynomial time.
3.2.2 Step 2.b: Construction of WAs/WTs that compute $f_{w,k},~{}g_{w,P}^{(1)},~{}g_{w,i,P}^{(2)}$
@@
The construction of $\mathcal{A}_{1}$ is relatively easy. As for $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$, the key observation stems from the Bayes’ formula:
$P(w|w\in L_{p})=\frac{P(w)\cdot I_{L_{p}}(w)}{P(L_{p})}$ (12)
[…]
We note that since $\mathcal{A}_{1},~{}\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ run in time polynomial in their respective input instances implies that the size of their output machines is also polynomial in the size of their input instance [^2] FP $\subset$ FPSPACE. This fact will appear explicitly in the constructive proof of lemma 3.5.
@@
The class of disjoint-DNFs (d-DNFs) is formally defined as follows:
Definition 4.1 (Disjoint DNF).
A d-DNF is a logical expression $\Phi(X_{1},X_{2},\ldots,X_{n})$ where $\{X_{1},X_{2},\ldots,X_{n}\}$ is a set of input boolean variables, such that:
  • $\Phi$ is expressed as a disjunction (logical OR) of clauses, where each clause is expressed as one or more conjunctions (logical AND) of literals.
  • Each clause in the expression is mutually exclusive from the others, ensuring that for any input combination $(X_{1},X_{2},\ldots,X_{n})$, only one clause evaluates to true.
@@
A Markovian distribution $P$ over a boolean random vector of dimension $N$ is given as:
@@
This result follows immediately from theorem 4.2, and the fact that given an arbitrary decision tree in DT, an equivalent d-DNF can be constructed in polynomial time with respect to the size of the decision tree (Property 1, (Aizenstein & Pitt, 1992)). ∎
@@ 5 Conclusion
Acknowledgements
We would like to express our gratitude to the anonymous reviewers for their invaluable feedback and constructive comments, which greatly contributed to enhancing the quality and clarity of this paper. In particular, we are thankful for their contributions in directing our attention to works addressing the limitations of the SHAP score.
Impact Statement
Overall, our research contributes to the burgeoning field of explainable artificial intelligence (XAI) by focusing on the SHAP score, a widely employed method for interpreting ML models. We believe that the newly introduced theoretical tools used in our work to substantiate our findings, particularly those that establish connections between boolean circuits and finite state automata, hold the potential to inspire the development of novel algorithms within the realm of formal XAI, thus contributing to advance the field as a whole. Ultimately, our work contributes to the ongoing efforts to promote transparency, accountability, and trustworthiness in AI systems, paving the way for their responsible deployment across various domains.
@@ Acknowledgements
$A_{w,P}$ refers to the WA that computes the language expressed in 19.
Given a string $w\in\Sigma^{*}$ and $P\in\texttt{MARKOV}$. A Markovian distribution over the finite support $\Sigma^{|w|}$ can be easily simulated by a DWA. The construction consists at maintaining in the state memory of the DWA the position reached so far in the sequence and the last generated symbol. These two pieces of information are sufficient to simulate a Markovian distribution.
@@
For an input instance $<\sigma^{\prime},n,\sigma,m,P>$, the quantity $\mathbf{G}(\sigma^{\prime},n,\sigma,m,P)$ refers to the conditional probability of generating a symbol in $L_{\sigma^{\prime}}$ at position $m$ given that the symbol $\sigma$ has been generated at position $n$. In essence, the computational problem $\mathbf{G}$ is reduced to the classical problem of inference in Bayesian Networks (Koller & Friedman, 2009). In general, the exact inference in Bayesian Networks is intractable (Tacettin & Ünlüyurt, 2005). However, in our case, leveraging the Markovian structure of the probability distribution enables building a tractable solution for the problem using a dynamic programming approach.